### v2

# The multiplier spectrum morphism is generically injective

Abstract.
In this paper, we consider the multiplier spectrum of periodic points, which is a natural morphism defined on the moduli space of rational maps on the projective line. A celebrated theorem of McMullen asserts that aside from the well-understood flexible Lattès family, the multiplier spectrum morphism is quasi-finite. In this paper, we strengthen McMullen’s theorem by showing that the multiplier spectrum morphism is generically injective. This answers a question of McMullen and Poonen.
1. Introduction
1.1. The multiplier spectrum morphism
For $d\geq 2,$ let ${\rm Rat}_{d}({\mathbb{C}})$ be the space of degree $d$ rational maps on ${\mathbb{P}}^{1}({\mathbb{C}})$. It is a smooth irreducible affine variety of dimension $12d+$ 1 [Sil12]. A rational map is called Lattès if it is semi-conjugate to an endomorphism on an elliptic curve. A Lattès map $f$ is called flexible Lattès if one can continuously vary the complex structure of the elliptic curve to get a family of Lattès maps passing through $f$. The structure of flexible Lattès maps is well-understood [Mil06, Lemma 5.5]. Let ${\rm FL}_{d}({\mathbb{C}})\subseteq{\rm Rat}_{d}({\mathbb{C}})$ be the locus of flexible Lattès maps, which is Zariski closed in ${\rm Rat}_{d}({\mathbb{C}})$. The group ${\rm PGL}\,_{2}({\mathbb{C}})={\rm Aut}({\mathbb{P}}^{1}({\mathbb{C}}))$ acts on ${\rm Rat}_{d}({\mathbb{C}})$ by conjugacy. The geometric quotient
${\mathcal{M}}_{d}({\mathbb{C}}):={\rm Rat}_{d}({\mathbb{C}})/{\rm PGL}\,_{2}({\mathbb{C}})$
is the (coarse) moduli space of endomorphisms of degree $d$ [Sil12]. The moduli space ${\mathcal{M}}_{d}({\mathbb{C}})={\rm Spec\,}({\mathcal{O}}({\rm Rat}_{d}({\mathbb{C}})))^{{\rm PGL}\,_{2}({\mathbb{C}})}$ is an irreducible affine variety of dimension $22d-$ 2 [Sil07, Theorem 4.36(c)], which is also a complex orbifold [Mil93], [Mil11]. Let $\Psi:{\rm Rat}_{d}({\mathbb{C}})\to{\mathcal{M}}_{d}({\mathbb{C}})$ be the quotient morphism. We set
${\mathcal{M}}_{d}^{*}({\mathbb{C}}):={\mathcal{M}}_{d}({\mathbb{C}})\setminus\Psi({\rm FL}_{d}({\mathbb{C}})).$
We note that ${\rm FL}_{d}({\mathbb{C}})=\emptyset$ when $d$ is not a square number, and $\Psi({\rm FL}_{d}({\mathbb{C}}))$ is an algebraic curve when $d$ is a square number.
As ${\mathcal{M}}_{d}$ is affine, to understand the geometry of ${\mathcal{M}}_{d}$, we only need to understand its function ring. There is a natural dynamically interesting family of morphisms from ${\mathcal{M}}_{d}({\mathbb{C}})$ to the affine spaces, which we call the multiplier spectrum morphisms. These morphisms give a natural system of functions on ${\mathcal{M}}_{d}({\mathbb{C}})$, which can be viewed as analogies of the theta functions on moduli space of elliptic curves (c.f. [Sil23, Remark 5.9]).
We now recall the construction. For every $f\in{\rm Rat}_{d}({\mathbb{C}})$ and $n\geq 1$, $f^{n}$ has exactly $N_{n}=d^{n}+1$ fixed points counted with multiplicity. The multiplier of a $f^{n}$-fixed point $x$ is the differential $df^{n}(x)\in{\mathbb{C}}$. Using elementary symmetric polynomials, their multipliers define a point $S_{n}(f)\in{\mathbb{C}}^{N_{n}}$. The multiplier spectrum of $f$ is the sequence $S_{n}(f),n\geq 1.$ Since $S_{n}$ take the same value in a conjugacy class of rational maps, each $S_{n}$ defines a morphism $S_{n}:{\mathcal{M}}_{d}({\mathbb{C}})\to{\mathbb{C}}^{N_{n}}$. Let $\tau_{d,n}$ be the morphism
$\tau_{d,n}:{\mathcal{M}}_{d}({\mathbb{C}})$ $\to{\mathbb{C}}^{N_{1}}\times\cdots\times{\mathbb{C}}^{N_{n}},$
$[f]$ $\mapsto(S_{1}(f),\dots,S_{n}(f)).$
Here for $f\in{\rm Rat}_{d}({\mathbb{C}})$, we define $\tau_{d,n}(f):=\tau_{d,n}([f])$.
For each $n\geq 1$, set
$R_{n}:=\{([f],[g])\in{\mathcal{M}}_{d}({\mathbb{C}})^{2}|\,\,\tau_{d,n}([f])=\tau_{d,n}([g])\}.$
Then $R_{n}$ are decreasing Zariski closed subsets of ${\mathcal{M}}_{d}({\mathbb{C}})^{2}$. By the Noetherianity, the sequence $R_{n}$ is stable for $n$ sufficiently large. Hence there exists a minimal positive integer $m_{d}$ such that $\tau_{d,m_{d}}(f)=\tau_{d,m_{d}}(g)$ implies that $\tau_{d,n}(f)=\tau_{d,n}(g)$ for every $n\geq 1$, i.e. $f$ and $g$ have the same multiplier spectrum. We define
$\tau_{d}:=\tau_{d,m_{d}}.$
It is well-known that elements in an irreducible component of ${\rm FL}_{d}({\mathbb{C}})$ have the same multiplier spectrum.
The following remarkable theorem of McMullen [McM87] says that outside ${\rm FL}_{d}({\mathbb{C}})$, the multiplier spectrum determines the conjugacy class of rational maps up to finitely many choices.
Theorem 1.1 (McMullen).
For every $d\geq 2$, the morphism
$\tau_{d}:{\mathcal{M}}_{d}^{*}({\mathbb{C}})\to{\mathbb{C}}^{N_{1}}\times\cdots\times{\mathbb{C}}^{N_{m_{d}}}$
is quasi-finite.
1.2. The multiplier spectrum morphism is generically injective
Definition 1.2.
For a point $x\in{\mathcal{M}}_{d}({\mathbb{C}})$, we say that $\tau_{d}$ is injective at $x$ if $\tau_{d}^{-1}(\tau_{d}(x))=\left\{x\right\}$. For a subset $X\subset{\mathcal{M}}_{d}({\mathbb{C}})$, we say that $\tau_{d}$ is injective on $X$ if $\tau_{d}$ is injective at every $x\in X$.
We quote the following question about the injectivity of $\tau_{d}$ from McMullen [McM87, Page 489]:
Noetherian properties imply there are an $N$ and an $M$ such that $E_{1}(R),\dots,E_{N}(R)$ determine $R$ up to at most $M$ choices, $\dots$ is $R$ determined uniquely? [^1] Following McMullen’s notations, $E_{i}(R)$ is the multiplier spectrum of periods $i$ of the rational map $R$.
It turns out that $\tau_{d}$ is not always injective on ${\mathcal{M}}_{d}^{*}({\mathbb{C}})$. Silverman showed that if $f$ is a rigid Lattès map defined over a number field $K$ whose class number is larger than 1, then $\tau_{d}$ is not injective at $[f]$ [Sil07, Theorem 6.62]. Another construction was introduced by Pakovich [Pak19a]. Let $f$ be a rational map. Following Pakovich [Pak19a], for any decomposition $f=h_{1}\circ h_{2}$ into a composition of rational maps, where $h_{1}$ and $h_{2}$ have degree at least 2, we say that the rational map $\tilde{f}:=h_{2}\circ h_{1}$ is an elementary transformation of $f$. We say that rational maps $f$ and $g$ are equivalent if there exists a chain of elementary transformations between $f$ and $g$. One can show that if $f$ and $g$ are equivalent then $\tau_{d}(f)=\tau_{d}(g)$, see Pakovich [Pak19b, Lemma 2.1]. Even though $\tau_{d}$ is not always injective on ${\mathcal{M}}_{d}^{*}({\mathbb{C}})$ as we have seen, one might hope that $\tau_{d}$ is injective at generic parameters, i.e. $\tau_{d}$ is injective on a Zariski open subset. By McMullen’s theorem (Theorem 1.1), there is an integer $m$, a non-empty Zariski open subset $U\subset{\mathcal{M}}_{d}({\mathbb{C}})$ and a Zariski open subset $W$ of the Zariski closure of $\tau_{d}(U)$ such that $\tau_{d}^{-1}(W)=U$ and $\tau_{d}:U\to W$ is a finite map of degree $m$. The integer $m$ is independent of the choices of $U$ and $W$, and we call this integer $m$ the degree of $\tau_{d}$. We denote this degree $m$ by $\deg(\tau_{d})$. When $\deg(\tau_{d})=1$, we say $\tau_{d}$ is generically injective.
Poonen asked whether $\tau_{d}$ is always generically injective [Sil12, Question 2.43]. The following is our main theorem. It gives an affirmative answer to Poonen’s question, hence an affirmative answer to McMullen’s question for generic parameters.
Theorem 1.3.
For every $d\geq 2$, the morphism
$\tau_{d}:{\mathcal{M}}_{d}({\mathbb{C}})\to{\mathbb{C}}^{N_{1}}\times\cdots\times{\mathbb{C}}^{N_{m_{d}}}$
is generically injective.
As a consequence, the system of functions given by the multiplier spectrum generate the fraction field of ${\mathcal{M}}_{d}.$
One can also consider the moduli space of polynomials and the multiplier spectrum morphism on it. For every $d\geq 2$ and $n\geq 1$, we let $\tilde{\tau}_{d,n}$ (resp. $\tilde{\tau}_{d}$) be the restriction of $\tau_{d,n}$ (resp. $\tau_{d}$) on the moduli space of polynomials. As the the moduli space of polynomials is a proper Zariski closed subset of ${\mathcal{M}}_{d}$, Theorem 1.3 does not implies the injectivity of $\tilde{\tau}_{d}$ directly. However, our proof of Theorem 1.3 still works in the polynomial setting.
Theorem 1.4.
For every $d\geq 2$, $\tilde{\tau}_{d}$ is generically injective.
The proof is of Theorem 1.4 is indeed even easier than the one for Theorem 1.3. To prove Theorem 1.3, a key ingredient is Pakovich’s result Theorem 3.3, which works only when $d\geq 4.$ So we need to apply Milnor [Mil93] and Gotou [Got23, Theorem 1.2] to treat the $d=2,3$ cases. To prove Theorem 1.4, we may replace Theorem 3.3 by [FG22, Theorem 23 and 24], which works for every $d\geq 2.$ We leave the proof of Theorem 1.4 for the readers.
While we were preparing this article, we knew from a private communication that Huguin have an independent proof of Theorem 1.4. Indeed Huguin proved the generic injectivity of $\tilde{\tau}_{d,2}$. From the communication with Huguin, we know that Huguin’s method is completely different from ours, which is based on Fujimura’s result [Fuj07] and on the computations by Gorbovickis [Gor15].
1.3. Ingredients in the proof of Theorem 1.3
To show Theorem 1.3, we argue by contradiction. Assume that $\tau_{d}$ is not generically injective. We first construct two algebraic families of rational maps $f_{t}$ and $g_{t}$, parametrized by the same algebraic curve $V$, such that: (1) for every $t\in V$, the images $f_{t}$ and $g_{t}$ in ${\mathcal{M}}_{d}$ are different; (2) $\tau_{d}(f_{t})=\tau_{d}(g_{t})$ for every $t\in V$. We can further ask that these two families satisfy several geometric and dynamical assumptions, c.f. Lemma 2.5. Our key step is to show that $f_{t}$ and $g_{t}$ are intertwined (c.f. Definition 3.1 ) for all but finitely many $t\in V$. The main ingredient of the proof for this step is the following variant of the Dynamical André-Oort (DAO) conjecture on the distribution of postcritically finite (PCF) maps.
Theorem 1.5.
Let $d\geq 2$ and let $f_{V},g_{V}$ be two degree $d$ non-isotrivial algebraic family parametrized by the same irreducible algebraic curve $V$, and $f_{V},g_{V}$ are not family of flexible Lattès maps. Assume that there are infinitely many $t\in V$ such that $f_{t}$ and $g_{t}$ are both PCF. Then for all but finitely many $t\in V$, $f_{t}$ and $g_{t}$ are intertwined.
From the above construction we can get a Zariski dense set of simple (c.f. Definition 3.2) rational maps $[f]$ in ${\mathcal{M}}_{d}({\mathbb{C}})$ which is intertwined with a simple rational map $g$ with $[g]\neq[f]$. We then use a result of Pakovich (c.f. Theorem 3.3) to get a contradiction.
We note that the DAO conjecture of Baker and DeMarco [BDM13] is considered an important conjecture in arithmetic dynamics, however not many of its applications are known. Theorem 1.3 is an application of the DAO conjecture.
In a private communication, DeMarco and Mavraki told us that when they preparing lectures in Harvard and Toronto in November 2023 [DM23], they find a simplification of our proof of Theorem 1.3. They observed that to prove Theorem 1.3, one may replace Theorem 1.5 by a weaker lemma (c.f. Lemma 3.8). This lemma can be proved following the same strategy of Theorem 1.5, but there are two important steps become easier. See Section 3.4 for details.
1.4. Previous results on the degree of the multiplier spectrum morphism
In [Gor15], Gorbovickis showed that $\tau_{d,n}$ is generically quasi-finite for $d\geq 2$ and $n\geq 3.$ A recursive formula for some upper bound of $\tau_{d,n}$ was obtained by Schmitt in the preprint version [Sch16]. An explicit upper bound of $\tau_{d,n}$ was obtained in Gotou’s recent work [Got23].
When $d=2$, Milnor [Mil93] showed that $\tau_{2,1}$ is in fact injective on ${\mathcal{M}}_{d}({\mathbb{C}})$ (see also [Sil12, Theorem 2.45]). In particular Theorem 1.3 holds when $d=2$. When $d=3$, Gotou showed that $\tau_{3,2}$ is generically injective (but not injective) [Got23, Theorem 1.2]. This was mentioned in [HT13a] which is the errata for [HT13b]. Previously there was no result about the generic injectivity of $\tau_{d}$ when $d\geq 4$.
One can also consider the moduli space of polynomials and the multiplier spectrum morphism on it. For every $d\geq 2$ and $n\geq 1$, we let $\tilde{\tau}_{d,n}$ be the restriction of $\tau_{d,n}$ on the moduli space of polynomials. In this case $\tilde{\tau}_{d,1}$ is generically quasi-finite (while $\tau_{d,1}$ is never generically quasi-finite, except when $d=2$). Fujimura showed that $\deg(\tilde{\tau}_{d,1})=(d-2)!$ [Fuj07]. The fiber structure of $\tilde{\tau}_{d,1}$ was studied by Sugiyama [Sug17], [Sug20], [Sug23].
1.5. Further results and problems
McMullen’s conjecture for the hyperbolicity of the moduli space ${\mathcal{M}}_{f}$
For a rational map $f\in{\rm Rat}_{d}({\mathbb{C}})$, McMullen and Sulllivan introduced the Teichmüller space ${\mathcal{T}}_{f}$ and the moduli space ${\mathcal{M}}_{f}$ [MS98]. We refer the readers to [MS98] and [Ast17] for the definitions of ${\mathcal{T}}_{f}$ and ${\mathcal{M}}_{f}$. The Teichmüller space ${\mathcal{T}}_{f}$ is a contractible complex manifold of dimension not larger than $22d-$ 2. The modular group ${\rm Mod}_{f}$ acts properly discontinuously on ${\mathcal{T}}_{f}$. The moduli space is the quotient ${\mathcal{M}}_{f}:={\mathcal{T}}_{f}/{\rm Mod}_{f}$. There is a natural holomorphic injection $\phi:{\mathcal{M}}_{f}\to{\mathcal{M}}_{d}({\mathbb{C}})$, such that $\phi({\mathcal{M}}_{f})$ is the set containing all quasiconformal conjugacy class of $f$.
In [McM87, Page 473], McMullen conjectured that for every $f\in{\rm Rat}_{d}({\mathbb{C}})$ that are not flexible Lattès, bounded holomorphic functions separate points on ${\mathcal{M}}_{f}$.
The above property of complex analytic spaces is sometimes called Carathéodory hyperbolic in the literature, which is a strong hyperbolic condition that implies Kobayashi hyperbolicity. In a forthcoming paper [JXar], we solve McMullen’s conjecture using the multiplier spectrum. We show a stronger result, namely for every $f\in{\rm Rat}_{d}({\mathbb{C}})$ that is not flexible Lattès, we show that there is a holomorphic injection $\tilde{\phi}:{\mathcal{M}}_{f}\to X_{f}$, where $X_{f}$ is a normal affine complex algebraic variety of dimension $22d-$ 2, such that $\tilde{\phi}({\mathcal{M}}_{f})$ is precompact in $X_{f}$. Moreover by applying Theorem 1.3, we get a explicit description of $X_{f}$ when $\dim{\mathcal{M}}_{f}=2d-2$ and $d\geq 4$.
Injective locus of $\tau_{d}$
We have seen before that there are two mechanisms that can produce rational maps with the same multiplier spectrum, one from Lattès maps, and another one from equivalent rational maps. Pakovich asked [Pak19a, Problem 3.1] whether these are the only obstructions of the injectivity of $\tau_{d}$. We conjecture that it is indeed the case.
Conjecture 1.6.
Let $f,g$ be rational maps of degree $d\geq 2$ such that the conjugacy classes of $f$ and $g$ are different. Assume that $\tau_{d}(f)=\tau_{d}(g)$, then one of the followings holds:
  • (i)
    $f$ and $g$ are Lattès maps;
  • (ii)
    $f$ is equivalent to $g$.
Generic injectivity of multiplier spectrum of small periods
In Theorem 1.1 and Theorem 1.3, the definition of $\tau_{d}$ requires the use of periodic points of periods not exceeding the number $m_{d}$, whose precise value is not effectively-known and is probably very large. It is interesting to know whether we can get generic injectivity using only periodic points of small periods. By dimension counting, $\tau_{d,1}$ is never generically injective except when $d=2$. However, we believe the following is true.
Conjecture 1.7.
For every $d\geq 2$, the morphism
$\tau_{d,2}:{\mathcal{M}}_{d}({\mathbb{C}})\to{\mathbb{C}}^{N_{1}}\times{\mathbb{C}}^{N_{2}}$
is generically injective.
Generic injectivity of the length spectrum
Replace the multipliers by their norm in the definition of multiplier spectrum, one gets the definition of the length spectrum. More precisely, for every $f\in\text{Rat}_{d}({\mathbb{C}})$ and $n\geq 1$, we denote by $L_{n}(f)\in{\mathbb{R}}_{\geq 0}^{N_{n}}$ the elements corresponding to the values of the elementary symmetric polynomials at the point $(|\lambda_{1}|,\dots,|\lambda_{N_{n}}|)\in{\mathbb{R}}_{\geq 0}^{N_{n}}$ where $\lambda_{i},i=1,\dots,N_{n}$ are the multipliers of all $f^{n}$-fixed points. The length spectrum of $f$ is defined by the sequence $L_{n}(f),n\geq 1.$ A priori, the length spectrum contains less information than the multiplier spectrum. In [JX23b], a parallel result of Theorem 1.1 has been shown, that is outside the flexible Lattès family, the length spectrum determines the conjugacy class of rational maps up to finitely many choices [JX23b, Theorem 1.5]. We believe the following parallel result of Theorem 1.3 for length spectrum is true. Note that ${\mathcal{M}}_{d}$ is defined over ${\mathbb{Q}}$ (hence over ${\mathbb{R}}$).
We propose the following conjecture as a parallel result of our Theorem 1.3.
Conjecture 1.8.
For every $d\geq 2$, there is a Zariski closed proper subset $E_{d}\subset{\mathcal{M}}_{d}$ defined over ${\mathbb{R}}$, such that for every $[f]\notin E_{d}$, if there is $g\in{\rm Rat}_{d}({\mathbb{C}})$ such that $f$ and $g$ have the same length spectrum, then $g$ or $\overline{g}$ (the complex conjugation of $g$) is contained in the conjugacy class of $f$.
More precisely, we propose the following description for rational maps with the same length spectrum.
Conjecture 1.9.
Let $f,g$ be non-Lattès rational maps of degree $d\geq 2$. If $f$ and $g$ has the same length spectrum, then $\tau_{d}(f)$ equals to either $\tau_{d}(g)$ or $\tau_{d}(\overline{g})$.
Multiplier spectrum for endomorphisms on ${\mathbb{P}}^{N}$
For holomorphic endomorphisms on ${\mathbb{P}}^{N}$, $N\geq 2$, one can also construct the corresponding moduli space ${\mathcal{M}}_{d}^{N}({\mathbb{C}})$ [Sil12], and the multiplier spectrum morphisms exist on ${\mathcal{M}}_{d}^{N}({\mathbb{C}})$ [Sil12]. Unlike in dimension one, essentially nothing is known about the multiplier spectrum morphisms when $N\geq 2$. It is of great interest to extend McMullen’s quasi-finiteness theorem (Theorem 1.1) and the generic injectivity theorem (Theorem 1.3) to higher dimension.
Acknowledgement
We thank Rin Gotou for sharing us his work on the generic injectivity of $\tau_{3,2}$ and let us know the result of Schmitt [Sch16]. We thank Fedor Pakovich, Valentin Huguin, Gabriel Vigny and Joseph Silverman for their comments and suggestions. We thanks Valentin Huguin for telling us his forthcoming proof for the generic injectivity of $\tilde{\tau}_{d,2}$. We thank Laura DeMarco and Niki Myrto Mavraki for sharing us their simplification of our proof of Theorem 1.3.
The first-named author would like to thank Beijing International Center for Mathematical Research in Peking University for the invitation. The second-named author Junyi Xie is supported by NSFC Grant (No.12271007).
2. Hyperbolic PCF maps of disjoint type
Definition 2.1.
A PCF map $f\in{\rm Rat}_{d}({\mathbb{C}})$ is called hyperbolic of disjoint type if $f$ has $22d-$ 2 number of distinct super-attracting cycles. Let $\underline{n}:=\left\{n_{1},\dots,n_{2d-2}\right\}\in({\mathbb{N}}^{*})^{2d-2}$, where ${\mathbb{N}}^{*}$ is the set of positive integers. A hyperbolic PCF map of disjoint type $f\in{\rm Rat}_{d}({\mathbb{C}})$ is called of type $\underline{n}$ if $f$ has $22d-$ 2 distinct super-attracting cycles of respective exact periods $n_{1},\dots,n_{2d-2}$.
For every $\underline{n}:=\left\{n_{1},\dots,n_{2d-2}\right\}\in({\mathbb{N}}^{*})^{2d-2}$, we let $X_{\underline{n}}\subset{\mathcal{M}}_{d}({\mathbb{C}})$ be the subset of all conjugacy classes $[f]$ such that $f$ is a hyperbolic PCF map of disjoint type $\underline{n}$. The following theorem was proved by Gauthier-Okuyama-Vigny in [GOV19, Theorem F].
Theorem 2.2.
For every sequence $\underline{n}(k)=(n_{1}(k),\dots,n_{2d-2}(k))\in({\mathbb{N}}^{*})^{2d-2}$ satisfying $\min_{1\leq j\leq 2d-2}n_{j}(k)\to+\infty$ when $k\to+\infty$, we have that $\cup_{k}X_{\underline{n}(k)}$ is Zariski dense in ${\mathcal{M}}_{d}({\mathbb{C}})$.
Definition 2.3.
Let $V$ be a quasi-projective variety over ${\mathbb{C}}$. An algebraic family on $V$ is an endomorphism $f_{V}$ on $V\times{\mathbb{P}}^{1}$ of the following form.
$f_{V}:V\times{\mathbb{P}}^{1}$ $\to V\times{\mathbb{P}}^{1}$
$(t,z)$ $\mapsto(t,f_{t}(z)).$
We say that $f_{V}$ has degree $d$ if for every $t\in V$, we have $\deg f_{t}=d.$ For a degree $d$ algebraic family $f_{V}$ on $V$, let $\Psi_{V}:V\to{\mathcal{M}}_{d}({\mathbb{C}})$ be the morphism sending $t\in V$ to the class of $f_{t}$ in ${\mathcal{M}}_{d}({\mathbb{C}}).$ We say that $f_{V}$ is isotrivial if $\Psi_{V}:V\to{\mathcal{M}}_{d}({\mathbb{C}})$ is locally constant.
The following lemma is a combination of [DF08, Proposition 2.4] and [DeM16, Theorem 1.1].
Lemma 2.4.
Let $f_{V}:V\times{\mathbb{P}}^{1}\to V\times{\mathbb{P}}^{1}$ be a degree $d$ non-isotrivial algebraic family, and let $a:V\to{\mathbb{P}}^{1}$ be a marked point. Assume that $a$ is not persistently preperiodic, then there exist infinitely many $t\in V$ such that $a(t)$ is periodic for $f_{t}$.
Proof.
By [DeM16, Theorem 1.1], the bifurcation locus $\text{Bif}(f_{V},a)$ is a non-empty closed set. Moreover $\text{Bif}(f_{V},a)$ is the support of a positive closed current with Hölder continuous potential [DS10, Lemma 1.1], hence $\text{Bif}(f_{V},a)$ has positive Hausdorff dimension [Sib99, Theorem 1.7.3], in particular $\text{Bif}(f_{V},a)$ is an infinite set. By definition, for every $t_{0}\in\text{Bif}(f_{V},a)$ and every open neighborhood $U$ of $t_{0}$, the family of maps $h_{n}:U\to{\mathbb{P}}^{1}$, $t\mapsto f^{n}_{t}(a(t))$ does not form a normal family. By [DF08, Proposition 2.4], there exists $t\in U$ such that $a(t)$ is periodic for $f_{t}$. Apply the above construction to infinitely many disjoint open subsets $U$ meeting $\text{Bif}(f_{V},a)$, we get infinitely many $t\in V$ such that $a(t)$ is periodic for $f_{t}$. ∎
For every $\underline{n}:=\left\{n_{1},\dots,n_{2d-3}\right\}\in({\mathbb{N}}^{*})^{2d-3}$ of $32d-$ 3 tuples, let $Y_{\underline{n}}\subset{\mathcal{M}}_{d}({\mathbb{C}})$ be the subset of all conjugacy classes $[f]$ such that $f$ has $32d-$ 3 distinct super-attracting cycles of respective exact periods $n_{1},\dots,n_{2d-3}$.
Lemma 2.5.
The following statements are true:
(1) The Zariski closure of $Y_{\underline{n}}$ is a (possibly reducible) algebraic curve provided that it is not empty and $Y_{\underline{n}}$ is Zariski open in its Zariski closure.
(2) The set $\cup_{\underline{n}}Y_{\underline{n}}$ is Zariski dense in ${\mathcal{M}}_{d}({\mathbb{C}})$.
(3) For every $\underline{n}$ such that $Y_{\underline{n}}$ is non-empty, and for every irreducible component $Y$ of $Y_{\underline{n}}$, there are infinitely many $[f]\in Y$ such that $f$ is a hyperbolic PCF map of disjoint type.
Proof.
We first show (1). Passing to a finite surjective morphism $\phi:V\to{\rm Rat}_{d}({\mathbb{C}})$, $t\mapsto f_{t}$, we can choose an algebraic family $f_{V}:V\times{\mathbb{P}}^{1}\to V\times{\mathbb{P}}^{1}$ such that all $22d-$ 2 critical points in this family can be algebraically parametrized by $c_{1},\dots,c_{2d-2}:V\to{\mathbb{P}}^{1}$. For every fixed $\underline{n}:=\left\{n_{1},\dots,n_{2d-3}\right\}\in({\mathbb{N}}^{*})^{2d-3}$, we define
$Z:=\left\{t\in V:f_{t}^{n_{j}}(c_{j})=c_{j},\;\;\text{for every}\;1\leq j\leq 2d-3\right\}.$
Then $Z$ is a Zariski closed subset in $V$. Since $\phi:V\to{\rm Rat}_{d}({\mathbb{C}})$ is surjective finite, and $\phi(Z)\subset{\rm Rat}_{d}({\mathbb{C}})$ is invariant by ${\rm PGL}\,_{2}({\mathbb{C}})$ conjugacy, $\Psi\circ\phi(Z)$ is Zariski closed in ${\mathcal{M}}_{d}({\mathbb{C}})$, where $\Psi:{\rm Rat}_{d}({\mathbb{C}})\to{\mathcal{M}}_{d}({\mathbb{C}})$ is the quotient morphism. Moreover by the definition of $Y_{\underline{n}}$, $Y_{\underline{n}}$ is Zariski open in $\Psi\circ\phi(Z)$. For each irreducible component $\tilde{Z}$ of $Z$, it is the intersection of at most $32d-$ 3 hypersurfaces, so $\tilde{Z}$ has codimension at most $32d-$ 3 in $V$. We claim that $\tilde{Z}$ has codimension $32d-$ 3 in $V$. Assume by contradiction it is not the case. Since critical orbit relations are constant along the fibers of the projection $\Psi\circ\phi:V\to{\mathcal{M}}_{d}({\mathbb{C}})$, $\Psi\circ\phi(\tilde{Z})$ has codimension $k<2d-3$ in ${\mathcal{M}}_{d}({\mathbb{C}})$. In particular the algebraic family $f_{\tilde{Z}}$ (the restriction of $f_{V}$ on $\tilde{Z}$) is non-isotrivial. Apply Lemma 2.4 to $f_{\tilde{Z}}$ and the marked point $c_{2d-2}:V\to{\mathbb{P}}^{1}$, there exists a positive integer $m$ such that the Zariski closed subset
$W:=\left\{t\in\tilde{Z}:f_{t}^{m}(c_{2d-2})=c_{2d-2}\right\}$
is non-empty. Moreover, $W$ has codimension at most $k+1<2d-2$. Then $\Psi\circ\phi(W)$ has codimension at most $k+1<2d-2$, which implies that $\Psi\circ\phi(W)$ contains a positive dimensional hyperbolic PCF family, this contradicts Thurston’s rigidity theorem [DH93] which says that the only positive dimensional PCF family in ${\mathcal{M}}_{d}({\mathbb{C}})$ is the flexible Lattès family. Hence $\tilde{Z}$ must have codimension $32d-$ 3 in $V$. This implies $\Psi\circ\phi(Z)$ has pure dimension 1. Then (1) holds since $Y_{\underline{n}}$ is Zariski open in $\Psi\circ\phi(Z)$.
Next we show (2). Since $\cup_{\underline{n}}X_{\underline{n}}\subset\cup_{\underline{n}}Y_{\underline{n}}$, by Lemma 2.2 (2), $\cup_{\underline{n}}Y_{\underline{n}}$ is Zariski dense.
Finally we show (3). By Lemma 2.5 (1) and by Lemma 2.4, there exist infinitely many $[f]$ in $Y_{\underline{n}}$ such that every citical point of $f$ is periodic and $f$ has $32d-$ 3 distinct super-attracting cycles of respective exact periods $n_{1},\dots,n_{2d-3}$. Since the set of conjugacy class $[f]$ such that all critical points of $f$ are periodic with bounded periods is finite, there exist infinitely many $[f]$ in $Y_{\underline{n}}$ such that $f$ has a periodic critical point with exact periods $N>\max_{1\leq j\leq 2d-3}n_{j}$, hence these $[f]$ are hyperbolic PCF maps of disjoint type. This implies (3). ∎
3. Generic injectivity of the multiplier spectrum morphism
3.1. Intertwined Rational maps
Definition 3.1.
Let $d\geq 2$ and $f,g\in{\rm Rat}_{d}({\mathbb{C}})$. We say $f$ and $g$ are intertwined if there exists a (maybe reducible) algebraic curve $Z\subset{\mathbb{P}}^{1}\times{\mathbb{P}}^{1}$ whose projections to both axis are onto, and $Z$ is invariant by the map $f\times g:{\mathbb{P}}^{1}\times{\mathbb{P}}^{1}\to{\mathbb{P}}^{1}\times{\mathbb{P}}^{1}$.
Definition 3.2.
A rational map $f\in{\rm Rat}_{d}({\mathbb{C}})$ is called simple if the $f$ has exactly $22d-$ 2 number of critical values.
Simple rational maps form a Zariski open subset of ${\rm Rat}_{d}({\mathbb{C}})$ for every $d\geq 2$.
The following theorem was obtained by Pakovich [Pak21] by combining several results in [Pak21]. We give a proof here for completeness.
Theorem 3.3.
Let $d\geq 4$. Then there exists a Zariski open subset $W_{d}$ of ${\rm Rat}_{d}({\mathbb{C}})$ such that :
  • (i)
    Elements in $W_{d}$ are simple;
  • (ii)
    Assume that $f$ and $g$ are intertwined, where $f\in W_{d}$ and $g$ is simple, then $f$ and $g$ are in the same conjugacy class.
Proof.
Let $d\geq 4$, by [Pak21, Theorem 1.2 and Lemma 3.7], there exists a Zariski open set $W_{d}\subset{\rm Rat}_{d}({\mathbb{C}})$ such that: (1) $W_{d}$ is contained in the set of simple rational maps, (2) if $f\in W_{d}$ and $f$ share the maximal entropy measure for some $g\in{\rm Rat}_{d}({\mathbb{C}})$, then $g=f$. We are going to show that the set $W_{d}$ satisfies Theorem 3.3 (ii).
Assume that $f\in W_{d}$, and $f$ is intertwined with a simple rational map $g\in{\rm Rat}_{d}({\mathbb{C}})$. By [Pak21, Theorem 1.4], there exist $m\geq 1$ and $\phi\in{\rm PGL}\,_{2}({\mathbb{C}})$ such that $(\phi g\phi^{-1})^{m}=f^{m}$. This implies $f$ and $\phi g\phi^{-1}$ have the same maximal entropy measure, hence we have $f=\phi g\phi^{-1}$. Hence $f$ and $g$ are in the same conjugacy class. ∎
3.2. A variant of the DAO conjecture
The following theorem is a variant of the main theorem in [JX23a]. In [JX23a, Theorem 1.2], the proof is for the product map $f_{V}\times f_{V}$ on ${\mathbb{P}}^{1}\times{\mathbb{P}}^{1}$, here we replace this product map by $f_{V}\times g_{V}$ on ${\mathbb{P}}^{1}\times{\mathbb{P}}^{1}$. The proof of Theorem 3.4 follows the same lines as previous proof of [JX23a, Theorem 1.3]. Indeed the proof of Theorem 3.4 is easier since we only need to construct a dynamical relation between $f_{V}$ and $g_{V}$.
Theorem 3.4 (=Theorem 1.5).
Let $d\geq 2$ and let $f_{V},g_{V}$ be two degree $d$ non-isotrivial algebraic family parametrized by the same irreducible algebraic curve $V$, and $f_{V},g_{V}$ are not family of flexible Lattès maps. Assume that there are infinitely many $t\in V$ such that $f_{t}$ and $g_{t}$ are both PCF. Then for all but finitely many $t\in V$, $f_{t}$ and $g_{t}$ are intertwined.
Proof.
Our proof is a modification of the one for [JX23a, Theorem 1.3].
We first reduce Theorem 3.4 to the case such that $V,f_{V},g_{V}$ are defined over $\overline{{\mathbb{Q}}}.$ To show this, let $\Psi_{f},\Psi_{g}:V\to{\mathcal{M}}_{d}$ be the morphisms sending $t$ to the conjugacy classes of $f_{t}$ and $g_{t}$ respectively. Denote by $\psi:V\to{\mathcal{M}}_{d}\times{\mathcal{M}}_{d}$ the morphism sending $t$ to $(\Psi_{f}(t),\Psi_{g}(t))$ and by $\Gamma$ the Zariski closure of its image. Since there are infinitely many $t\in V$ such that $f_{t}$ and $g_{t}$ are both PCF and $f_{V},g_{V}$ are not family of flexible Lattès maps, by Thurston’s rigidity theorem for PCF maps [DH93], $\Gamma$ is defined over $\overline{{\mathbb{Q}}}.$ Consider the natural morphism $\Psi\times\Psi:{\rm Rat}_{d}\times{\rm Rat}_{d}\to{\mathcal{M}}_{d}\times{\mathcal{M}}_{d}$. There is an algebraic curve $V^{\prime}$ in $(\Psi\times\Psi)^{-1}(\Gamma)$ defined over $\overline{{\mathbb{Q}}}$ such that $\Psi\times\Psi(V^{\prime})$ is dense in $\Gamma.$ Then $V^{\prime}$ defines algebraic families $f^{\prime}_{V^{\prime}},g^{\prime}_{V^{\prime}}$ of degree $d$ maps parametrized by $V^{\prime}$. Both of them are defined over $\overline{{\mathbb{Q}}}.$ To prove Theorem 3.4 for $f_{V},g_{V}$, we only need to prove it for $f^{\prime}_{V^{\prime}},g^{\prime}_{V^{\prime}}$. So we may assume now that $V,f_{V},g_{V}$ are defined over $\overline{{\mathbb{Q}}}.$
After replacing $V$ by its normalization and then by some finite ramification cover, we may assume that $V$ is smooth and both $f_{V}$ and $g_{V}$ have exactly $22d-$ 2 marked critical points $a_{1},\dots,a_{2d-2}$ and $b_{1},\dots,b_{2d-2}.$
Now we follow the notations in [JX23a]. We denote by $\mu_{f,a_{i}},\mu_{g,b_{i}}$ the bifucation measures on $V({\mathbb{C}})$ for pairs $(f_{V},a_{i})$ and $(g_{V},b_{i})$ respectively. By [DF08, Theorem 2.5] (and DeMarco [DeM16]), $\mu_{f_{V},a_{i}}$ (resp. $\mu_{g_{V},b_{i}}$) is vanishes if and only if $a_{i}$ (resp. $b_{i}$) is preperiodic. In this case, they are called passive. Otherwise, they are called active. Let $\mu_{f_{V},{\rm bif}}:=\sum_{i=1}^{2d-2}\mu_{f_{V},a_{i}}$ and $\mu_{g_{V},{\rm bif}}:=\sum_{i=1}^{2d-2}\mu_{g_{V},b_{i}}$ be the bifurcation measures for $f_{V}$ and $g_{V}$ respectively. By Thurston’s rigidity theorem for PCF maps [DH93], both $\mu_{f_{V},{\rm bif}}$ and $\mu_{g_{V},{\rm bif}}$ are non-zero. We may assume that $a_{1}$ and $b_{1}$ are active.
By [JX23a, Corollary 2.4], for every active $a_{i}$ (resp. $b_{i}$), $\mu_{f_{V},a_{i}}$ (resp. $\mu_{g_{V},b_{i}}$) and $\mu_{f_{V},{\rm bif}}$ (resp. $\mu_{g_{V},{\rm bif}}$) are proportional. Moreover the proof of [JX23a, Corollary 2.4] implies that $\mu_{f_{V},{\rm bif}}$ and $\mu_{g_{V},{\rm bif}}$ are proportional. Let $\mu$ be the probability measure on $V({\mathbb{C}})$ which is proportional to $\mu_{f_{V},{\rm bif}}$ (hence $\mu_{g_{V},{\rm bif}}$).
Since aside from the flexible Lattès locus, the exceptional maps  [^2] As in [JX23b, Section 1.1], we call $g$ exceptional if it is a Lattès map or semiconjugates to a monomial map. are isolated in ${\mathcal{M}}_{d}({\mathbb{C}})$. As $\mu$ has continuous potential, we have the following property:
  • (1)
    For $\mu$-a.e. point $t\in V({\mathbb{C}})$, both $f_{t}$ and $g_{t}$ are non-exceptional.
Let ${\rm Corr}({\mathbb{P}}^{1})^{f_{t}\times g_{t}}_{*}$ be the set of $f_{t}\times g_{t}$-invariant Zariski closed subsets $\Gamma_{t}\subseteq{\mathbb{P}}^{1}\times{\mathbb{P}}^{1}$ of pure dimension $1$ such that both $\pi_{1}|_{\Gamma_{t}}$ and $\pi_{2}|_{\Gamma_{t}}$ are finite, where $\pi_{1},\pi_{2}$ are the first and the second projections. Let ${\rm Corr}^{\flat}({\mathbb{P}}^{1}_{V})^{f_{V}\times_{V}g_{V}}_{*}$ be the set of $f_{V}\times_{V}g_{V}$-invariant Zariski closed subsets $\Gamma\subseteq V\times({\mathbb{P}}^{1}\times{\mathbb{P}}^{1})$ which is flat over $V$ and whose generic fiber is in ${\rm Corr}({\mathbb{P}}^{1}_{\eta})^{f_{\eta}\times g_{\eta}}_{*}$, where $\eta$ is the generic point of $V$. In general, a correspondence $\Gamma_{t}\in{\rm Corr}({\mathbb{P}}^{1})^{f_{t}\times g_{t}}_{*}$ may not be contained in any correspondence in ${\rm Corr}^{\flat}({\mathbb{P}}^{1}_{V})^{f_{V}\times_{V}g_{V}}_{*}$. On the other hand, by the proof of [JX23a, Proposition 3.11], it is the case if $t$ is transcendental i.e. $t\in V({\mathbb{C}})\setminus V(\overline{{\mathbb{Q}}})$ As $\mu$ has continuous potential, $\mu$-a.e. point $t\in V({\mathbb{C}})$ is transcendental. We then get the following property:
  • (2)
    For $\mu$-a.e. point $t\in V({\mathbb{C}})$, every $\Gamma_{t}\in{\rm Corr}({\mathbb{P}}^{1})^{f_{t}\times g_{t}}_{*}$ is contained in some correspondence in ${\rm Corr}^{\flat}({\mathbb{P}}^{1}_{V})^{f_{V}\times_{V}g_{V}}_{*}$.
Next, we consider some typical non-uniformly hyperbolic conditions. See [JX23a, Definition 4.1, 5.1 and 8.1] for the definitions of such conditions. By [JX23a, Proposition 8.3], there exists $\lambda_{0}>1$ such that
  • (3)
    for $\mu-$a.e. point $t$, $f_{t},g_{t}$ are ${\rm CE}(\lambda_{0})$ hence ${\rm TCE}(\lambda_{1})$ for every $1<\lambda_{1}<\lambda_{0}$, by Przytycki-Rohde [PR98].
By [JX23a, Theorem 4.3], which is essentially due to De Thélin-Gauthier-Vigny [DTGV21],
  • (4)
    for $\mu-$a.e. point $t$, $f_{t},g_{t}$ are ${\rm PCE}(\lambda)$ for some $1<\lambda_{2}<d^{1/2}$.
By [JX23a, Theorem 4.6],
  • (5)
    for $\mu-$a.e. point $t$, $f_{t},g_{t}$ are ${\rm PR}(s)$ for some $s>1/2$.
As we have the conditions (1), (3), (4), (5), we may apply the proof of [JX23a, Proposition 7.8] for the active marked points $a_{1},b_{2}$ to show that, for $\mu$-a.e. point $t\in V({\mathbb{C}})$, there is $\Gamma_{t}\in{\rm Corr}({\mathbb{P}}^{1})^{f_{t}\times g_{t}}_{*}$. By condition (2), ${\rm Corr}^{\flat}({\mathbb{P}}^{1}_{V})^{f_{V}\times_{V}g_{V}}_{*}\neq\emptyset$. This implies that for all but finitely many $t\in V$, $f_{t}$ and $g_{t}$ are intertwined, which concludes the proof. ∎
3.3. Proof of the generic injectivity
Proof of Theorem 1.3.
The case $d=2$ was proved by Milnor [Mil93] (see also [Sil12, Theorem 2.45]). The case $d=3$ was proved by Gotou [Got23, Theorem 1.2] (ee also [HT13a] which is the errata for [HT13b]).
Now we assume that $d\geq 4$. Assume by contradiction that $\tau_{d}$ is not generically injective. By Theorem 3.3, there is a non-empty Zariski open subset $U$ of ${\mathcal{M}}_{d}^{*}$ such that for every $f\in{\rm Rat}_{d}({\mathbb{C}})$ with $[f]\in U$, $f$ satisfies the two conditions in Theorem 3.3. There is a Zariski open subset $W$ of the Zariski closure of $\tau_{d}(U)$ such that $\tau_{d}^{-1}(W)\subseteq U$ and $\tau_{d}|_{\tau_{d}^{-1}(W)}:\tau_{d}^{-1}(W)\to W$ if finite étale of degree at least $2$. After shrinking $U$, we may assume that $U=\tau_{d}^{-1}(W).$ We fix this Zariski open subset $U$.
For an algebraic family of rational maps $f$, we let $\Psi_{f}:V\to{\mathcal{M}}_{d}$ be the morphism sending $t$ to the conjugacy class of $f_{t}$. Now we construct two non-isotrivial algebraic families $f_{V},g_{V}$ of degree $d$ rational maps parametrized by the same irreducible algebraic curve $V$ such that the following holds: There exists $\underline{n}:=\left\{n_{1},\dots,n_{2d-3}\right\}\in({\mathbb{N}}^{*})^{2d-3}$ such that we have
(3.1) $\Psi_{f}(V)\subset Y_{\underline{n}}\cap U,\;\text{and}\;\Psi_{g}(V)\subset U,$
where $Y_{\underline{n}}$ is the algebraic curve in Lemma 2.5, and we have
(3.2) $\tau_{d}\circ\Psi_{f}=\tau_{d}\circ\Psi_{g},$
finally for every $t\in V$, we have
(3.3) $\Psi_{f}(t)\neq\Psi_{g}(t).$
We now describe this construction. By Lemma 2.5 (2), there exists $\underline{n}:=\left\{n_{1},\dots,n_{2d-3}\right\}\in({\mathbb{N}}^{*})^{2d-3}$ such that $Y_{\underline{n}}\cap U\neq\emptyset.$ We pick an irreducible component $Z$ of $Y_{\underline{n}}$ meeting $U.$ By Lemma 2.5 (1), $Z$ is an algebraic curve. Set $Z_{1}:=Z\cap U.$ Then $Z^{\prime}_{1}:=\tau_{d}(Z_{1})$ is an algebraic curve. Consider the fiber product $X:=Z_{1}\times_{Z^{\prime}_{1}}(\tau_{d}|_{U})^{-1}(Z^{\prime}_{1}).$ Recall that by the construction of $U$, $\tau_{d}|_{U}:U\to W$ is finite étale of degree at least $2$. This implies that there is an irreducible component $\tilde{Z}$ of $X$ which is not the diagonal. We denote by $\pi_{1}:X\to Z_{1}$ and $\pi_{2}:X\to(\tau_{d}|_{U})^{-1}(Z^{\prime}_{1})$ for the first and the second projections. Both of $\pi_{1}$ and $\pi_{2}$ are finite étale. Set $Z_{2}:=\pi_{2}(\tilde{Z}).$ For every $z\in\tilde{Z}$, $\pi_{1}(z)\neq\pi_{2}(z).$
Pick an irreducible curve $C_{1}^{\prime}$ in $\Psi^{-1}(Z_{1})\subseteq{\rm Rat}_{d}({\mathbb{C}})$, such that $\Psi|_{C_{1}^{\prime}}:C_{1}^{\prime}\to Z_{1}$ is dominant, where $\Psi:{\rm Rat}_{d}({\mathbb{C}})\to{\mathcal{M}}_{d}({\mathbb{C}})$ is the quotient morphism. Pick an irreducible component $C_{1}$ of $C_{1}^{\prime}\times_{Z_{1}}\tilde{Z}$. Let $\phi_{1},\phi_{2}$ be the projection to the first and the second coordinates. We then get an algebraic family $h_{1}$ of degree $d$ rational maps parametrized by $C_{1}$ such that for every $t\in C_{1}$, the map $f_{t}$ is the one given by $\phi_{1}(t)\in{\rm Rat}_{d}({\mathbb{C}})$. Then the morphism $\Psi_{1}:C_{1}\to{\mathcal{M}}_{d}({\mathbb{C}})$ sending $t$ to the conjugacy class of $f_{t}$ is the composition of $\phi_{2}:C_{1}\to\tilde{Z}$ and $\pi_{1}:\tilde{Z}\to Z_{1}\subseteq{\mathcal{M}}_{d}({\mathbb{C}})$.
Similarly, we get an algebraic family $h_{2}$ of degree $d$ rational maps parametrized by an irreducible curve $C_{2}$ and a dominant morphism $\phi_{2}^{\prime}:C_{2}\to Z_{2}$ such that $\Psi_{2}:C_{2}\to{\mathcal{M}}_{d}({\mathbb{C}})$ sending $t$ to the conjugacy class of $f_{t}$ is the composition of $\phi_{2}^{\prime}:C_{2}\to\tilde{Z}$ and $\pi_{2}:\tilde{Z}\to Z_{2}\subseteq{\mathcal{M}}_{d}({\mathbb{C}}).$
Pick an irreducible component $V$ of $C_{1}\times_{\tilde{Z}}C_{2}$ and denote by $\beta_{i},i=1,2$ the projections to the first and the second coordinates. They induce two algebraic families $f_{V},g_{V}$ of degree $d$ maps parametrized by the same irreducible algebraic curve $V$. The image of the morphism $\Psi_{f}:V\to{\mathcal{M}}_{d}({\mathbb{C}})$ sending $t$ to the conjugacy class of $f_{t}$ is Zariski dense in $Z_{1}$ and the image of $\Psi_{g}:V\to{\mathcal{M}}_{d}({\mathbb{C}})$ sending $t$ to the conjugacy class of $g_{t}$ is Zariski dense in $Z_{2}.$ Hence $f_{V}$ and $g_{V}$ are non-isotrivial. By our construction (3.1) automatically holds. Moreover we have
$\tau_{d}\circ\Psi_{f}=\tau_{d}\circ\Psi_{g}.$
As for every $z\in\tilde{Z}$, $\pi_{1}(z)\neq\pi_{2}(z)$, for every $t\in V$, we have
$\Psi_{f}(t)\neq\Psi_{g}(t).$
Thus we complete the construction.
We need the following lemma.
Lemma 3.5.
Let $d\geq 2$ and let $f,g\in{\rm Rat}_{d}({\mathbb{C}})$ have the same multiplier spectrum. If $f$ is a hyperbolic PCF map of disjoint type $\underline{n}:=\left\{n_{1},\dots,n_{2d-2}\right\}\in({\mathbb{N}}^{*})^{2d-2}$, then $g$ is also a hyperbolic PCF map of disjoint type $\underline{n}$.
Proof.
We will show that for every $f\in{\rm Rat}_{d}({\mathbb{C}})$, we can read the number of the critical cycles and length of them. Clearly, this implies Lemma 3.5.
For $n\geq 1$, we view $S_{n}(f)$ as the multi-set $\{df^{n}(x)|\,\,x\in\mathrm{Fix}(f^{n})\}$. It has exactly $d^{n}+1$ elements. We denote by $u_{n}$ the number of zeros in $S_{n}(f)$. It is clear that the sequence $u_{n}$ is determined by the multiplier spectrum of $f$. For every $l\geq 1$, denote by $m_{l}$ the number of critical cycles of length $l$. It is clear that the sequence $m_{l},l\geq 1$ determines the number of the critical cycles and length of them. So we only need to read the sequence $m_{l},l\geq 1$ from $u_{n},\geq 1.$ It is clear that $u_{n}=\sum_{l|n}m_{l}.$ By Mobius inversion formula, we get $m_{n}=\sum_{l|n}\mu(l)u_{n/l}$, where $\mu:{\mathbb{N}}^{*}\to\{-1,0,1\}$ is the Möbius function. This concludes the proof. ∎
We continue the proof of Theorem 1.3. By Lemma 2.5 (3), there are infinitely many $t\in V$ such that $f_{t}$ is a hyperbolic PCF map of disjoint type. By (3.2), $f_{t}$ and $g_{t}$ have the same multiplier spectrum, hence by Lemma 3.5, for such $t$, $g_{t}$ is also a hyperbolic PCF map of disjoint type. By Theorem 3.4, after shrinking $V$, $f_{t}$ and $g_{t}$ are intertwined for every $t\in V$. Since (3.1) holds and $f_{t}$ and $g_{t}$ are intertwined for every $t\in V$, by Theorem 3.3, $\Psi_{f}(t)=\Psi_{g}(t)$ for every $t\in V$. This contradicts (3.3). We then conclude the proof. ∎
Remark 3.6.
In the proof Theorem 1.3, we do not need the full strength of Lemma 3.5. One can replace it with the following fact: if two rational maps $f$ and $g$ have the same multiplier spectrum such that $f$ is PCF, then $g$ is also PCF. This is a consequence of [JXZ23, Theorem 1.12].
Remark 3.7.
There is also a way to prove Theorem 1.3 considering all PCF maps instead of hyperbolic PCF maps of disjoint type. In such a way, we may replace Gauthier-Okuyama-Vigny’s result [GOV19, Theorem F] (c.f. Theorem 2.2) by the well-known fact that the PCF parameters are Zariski dense in ${\mathcal{M}}_{d}$. On the other hand, Lemma 3.5 is not sufficient for the proof and we need to use [JXZ23, Theorem 1.12] instead. As [JXZ23, Theorem 1.12] relies on Siegel’s deep theorem for integer points and the proof of Lemma 3.5 is elementary, we choose the current proof.
3.4. An alternative approach
In a private communication, DeMarco and Mavraki told us that when they preparing lectures in Harvard and Toronto in November 2023 [DM23], they find a simplification of our proof of Theorem 1.3.
Their key observation is as follows: A key ingredient of Theorem 1.3 is Theorem 3.4 (=Theorem 1.5) which proves more than what we need. The proof of Theorem 3.4 strongly relies on the results and methods developed in author’s previous paper [JX23a].
First, in the proof of Theorem 1.3, we do not need Theorem 3.4 for general families. If we choose the way to prove Theorem 1.3 as in Remark 3.7, we only need to apply Theorem 3.4 for families $f_{V},g_{V}$ having the following additional assumptions:
  • (A)
    For every $t\in V({\mathbb{C}})$, the multiplier spectrums for $f_{t}$ and $g_{t}$ are the same.
  • (B)
    The curve $V$ is smooth and both $f_{V}$ and $g_{V}$ have exactly $22d-$ 2 marked critical points $a_{1},\dots,a_{2d-2}$ and $b_{1},\dots,b_{2d-2}$. Moreover all these marked critical points except $a_{1},b_{1}$ are strictly preperiodic at every $t\in V$.
Second, in Theorem 3.4, we proved that $f_{t}$ and $g_{t}$ are intertwined for all but finitely many $t\in V$. However, to prove Theorem 1.3, we only need one such $t.$ So we may replace Theorem 3.4 to the following weaker result.
Lemma 3.8.
Let $f_{V},g_{V}$ be families as in Theorem 3.4. We further assume (A), (B) holds. Then there is $t_{0}\in V({\mathbb{C}})$ such that $f_{t_{0}}$ and $g_{t_{0}}$ are intertwined.
To prove Lemma 3.8, one can follow the same strategy of the proof of Theorem 3.4 (hence the proof of [JX23a, Theorem 1.3]). But there are two important steps become easier. After replacing $V$ by its normalization and then by some finite ramification cover, we may assume that $V$ is smooth and both $f_{V}$ and $g_{V}$ have exactly $22d-$ 2 marked critical points $a_{1},\dots,a_{2d-2}$ and $b_{1},\dots,b_{2d-2}.$
In the first step of the proof of Theorem 3.4, we show that there is a probability measure on $V({\mathbb{C}})$ which is proportional to $\mu_{f_{V},a_{i}}$ (resp. $\mu_{g_{V},b_{i}}$) for every active $a_{i}$ (resp. $b_{i}$). The proof is based on Yuan-Zhang’s deep equidistribution theorem [YZ21, Theorem 6.2.3] on quasi-projective varieties. The corresponding step for Lemma 3.8 is much easier. By assumption (B), $a_{1},b_{1}$ are the only active marked points for $f_{V}$ and $g_{V}$ respectively. Hence $\mu_{f_{V},a_{1}}$ (resp. $\mu_{g_{V},b_{1}}$) is proportional to $\mu_{f_{V},{\rm bif}}$ (resp. $\mu_{g_{V},{\rm bif}}$). By [Oku15] the Lyapunov exponent $L(h)$ for every rational function $h$ on ${\mathbb{P}}^{1}({\mathbb{C}})$ can be computed using its mutiplier spectrum. Then (A) implies that $L(f_{t})=L(g_{t})$ for every $t\in V({\mathbb{C}}).$ By [DeM01, DeM03], the function $t\in V({\mathbb{C}})\mapsto L(f_{t})$ (resp. $t\in V({\mathbb{C}})\mapsto L(g_{t})$) is a potential of $\mu_{f_{V},{\rm bif}}$ (resp. $\mu_{g_{V},{\rm bif}}$). This gives a logically easier proof of this step for Lemma 3.8.
Another crucial step to prove Theorem 3.4 is to construct similarities between the bifurcation measure on the parameter space $V({\mathbb{C}})$ and the maximal entropy measures on the phase spaces. As the conclusion of Theorem 3.4 is for all but finitely many parameters in $V({\mathbb{C}})$, we construct such similarities for $\mu$-a.e. $t\in V({\mathbb{C}})$. This step is essentially done in [JX23a], which is technically difficult. We first show that for a $\mu$-generic parameter $t$, $f_{t}$ and $g_{t}$ satisfy some Collet-Eckmann-type conditions (which can be thought as some weak hyperbolicity). We then built the similarities under such conditions. As our hyperbolicity assumption is very weak, we can not apply the previous methods as in [Tan90, FG22, Gau22] to construct similarities. Actually, our proof is based on subtle estimates of distortions for non-injective maps and a suitable binding argument. This step becomes much easier for Lemma 3.8, as we only need to construct the similarities at a single parameter $t_{0}$. By [Duj13, Theorem 0.1], there is a dense set $P$ of parameters $t\in{\rm Supp\;}\,\mu_{f_{V},a_{1}}$ such that $a_{1}$ is transversally pre-repelling at $t$. By (B), for every $t\in P$, $f_{t}$ is PCF with no periodic critical point. The proof of Lemma 3.5 implies that for every $t\in P$, $g_{t}$ is also PCF with no critical periodic point. In particular, $b_{1}(t)$ is pre-repelling for $g_{t}$. Pick any point $t_{0}\in P$, it is easy to get the similarities at $t_{0}$ applying the method in [FG22, Gau22].